. 40% of the population has type A blood. a. If 9 people are selected at random, what is the probability that less than four of them have type A blood. b. If 90 donors come to give blood one day, what is the probability that less than 40 of them have Type A blood (using the normal approximation)? Explain why this is higher or lower than the answer in part (a). c. If 15 people give blood one day, what is the probability that more than 10 of them are Type A? d. If 150 people come to give blood, what is the probability that more than 100 the donors is of Type A? Explain why this answer is higher or lower than the answer in part (c).
Answer #1

P(Type A) = 0.4

a.

P(x<4) = 0.4826

b.

Mean = n*p = 90*0.4 = 36

Standard dev = (n*p*q)^0.5 = (90*0.4*0.6)^0.5 = 4.6476

c.

P(x>10) = 0.0093

d.

Mena = n*p = 150*0.4 = 60

Standard dev = (n*p*q)^0.5 = 6

P(X>100)